OpenStudy (anonymous):

Using complete sentences, explain how you would factor completely x^9 –27

OpenStudy (cwrw238):

you can use the identity a^3 - b^3 = (a - b)(a^ - can't remember it - ill recheck

jimthompson5910 (jim_thompson5910):

a^3 - b^3 = (a-b)(a^2 + ab + b^2) is what you're looking for cwrw238

OpenStudy (cwrw238):

thats it - thanx lol

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (cwrw238):

so put a = x^3 and b = 3 into Jim's formula

OpenStudy (cwrw238):

(x^3)^3 - 3^3 = (x^3 - 3)(x^3)^2 + 3x^3 + 3^2) = (x^3 - 3)(x^6 + 3x^3 + 9)

OpenStudy (anonymous):

x^9 –27=0 ...firstly u take 27 to the right and u will get x^9 =27 then apply the 9th root both sides u will get x=\[\sqrt[9]{27}\] then x = (27)^(1/9) x=(3^3)^(1/9) x=3^(1/3) =\[\sqrt[3]{3}\] ........so ....u can use a calculator 2 find the answer if uwant 2 go further .....but it is not important